# EnsAI: An Emulator for Atmospheric Chemical Ensembles

### v1

@@ 8.4 Comparison Results for Spatial Correlations
To conclude this section, we briefly examine the higher correlation length modes with $m>2$. Figure 13 shows the $RMSE$ and $ACC$ values of $L_{ec}[m]/L_{ec}[0]$, taken over the evaluation time period and model domain, for $m$ between 1 and 15. From the top plot, we can see that the $RMSE$ values for both the EnsAI ensemble and static covariance decrease with increasing wavenumber, as all three covariances typically have less features on smaller scales (the decreasing values of $L_{ec}[m]$ with increasing $m$ that was shown in the example in Fig. 11 is a typical feature of the horizontal correlations). The $RMSE$ values for the EnsAI ensemble are smaller than those for the static covariance for all wavenumbers, but the difference between the two decrease with increasing values of $m$ and nearly converge by $m=15$. As seen in bottom panel of Fig. 13, the $ACC$ values for the EnsAI ensemble is above 0.85 for $m=1$ and decreases with increasing $m$, but is still at a value just below 0.5 by $m=15$. In contrast, the $ACC$ values for the static covariance show little change with wavenumber and is below 0.1 for all wavenumbers.
Figure 13: $RMSE$ and $ACC$ values for the horizontal correlation length modes for the emissions/surface concentration correlation $L_{ec}[m]$ as a function of wavenumber $m$ for the EnsAI ensemble (orange) and the static covariance (green).
@@ 9 Inversions
Sections 8.3 and 8.4 showed the results of direct comparisons between the covariances to determine how well EnsAI replicated the GEM-MACH ensemble perturbations. For some applications, such as model uncertainty quantification, the determination of the model’s error covariance is the end goal. However, for inversions/data assimilation, the objective is to produce an analysis (as well as forecasts initialized by the analysis in the case of data assimilation). In this section, we compare emission inversion results made using the different model error covariances.
@@
As the inversions are only sensitive to the model error covariances at times and locations near the available observations, not all differences between the error covariances will show in the comparison between the increments and will be dependent on the particular observations used. As such, pairing the comparison of inversion results with the direct comparison between the covariances in the previous two sections provides a robust examination between the different error covariances.
The ammonia emission increments produced using the inversion scheme described in Section 2 for the two test time periods are displayed in Figure 14. The left column shows the emission increments produced by inversions that used the original GEM-MACH ensemble. As found in Sitwell et al. (2022), the inversions increase the ammonia emissions in most parts of the model domain, with a few decreases in June in some parts of North Carolina and in the midwestern United States. We note that the emissions increments are smaller than those computed in Sitwell et al. (2022), since the inversions in this work each only use a week’s worth of observations instead of a full month’s worth. However, since our goal is only to determine how well inversions that use EnsAI reproduces the increments made using the original GEM-MACH ensemble, this should not be relevant for our comparisons.
The middle column in Fig. 14 shows the difference between the increments produced using the EnsAI ensemble and those produced using the GEM-MACH ensemble (as displayed in the left column), while the right column shows the difference between the increments made using the static covariance and those made using the GEM-MACH ensemble. Looking at the middle column of Fig. 14, the EnsAI and GEM-MACH increments are mostly within 5 $\rm{mg}\;\rm{s}^{-1}\;\rm{km}^{-2}$ of one another, corresponding to being within roughly 15% of one another. The only major exception to this is in north-wester Iowa in the June inversion, where the EnsAI increment is only around -12 $\rm{mg}\;\rm{s}^{-1}\;\rm{km}^{-2}$ instead of about -25 $\rm{mg}\;\rm{s}^{-1}\;\rm{km}^{-2}$ for the GEM-MACH inversion. By contrast, the right column of Fig. 14 shows that there are a number of locations (North Carolina, Pennsylvania, southern Nebraska, central Texas) where the static covariance and GEM-MACH increments differ by over 15 $\rm{mg}\;\rm{s}^{-1}\;\rm{km}^{-2}$. The increments produced using the static covariances differs from the GEM-MACH increment by a larger amount than the EnsAI increment does in most locations within the model domain.
Figure 14: Ammonia emissions increments for inversions using CrIS ammonia retrievals for the week of 2016-06-01 to 2016-06-07 (top row) and for the week of 2016-08-01 to 2016-08-07 (bottom row). The left column shows the increments produced with inversions that use the GEM-MACH ensemble. The middle column shows the difference between increments produced using the EnsAI ensemble and increments produced using the GEM-MACH ensemble, while the right column shows the difference between the static error covariance increments and the GEM-MACH ensemble increments.
@@ 10 Conclusions
Many geophysical AI-based models use extensive training data sets, such as the ERA5 or CAMS analysis/reanalysis. However, EnsAI was trained on a much smaller data set that was only comprised of 15 weeks of data. This suggests that ensemble generation systems that are only designed to predict model perturbations (instead of the model trajectory itself) seems to produce good results using much smaller training data sets.
@@
While application to emissions inversion systems is a substantial development by itself, the longer term goal is expand EnsAI to more atmospheric chemical constituents to use in chemical data assimilation. Although EnsAI was able to reproduce most the time-dependent features of the ammonia ensemble with only surface fields as input, more vertical levels may be needed if applied to other species with longer atmospheric life times.
@@ Appendix A Ensemble-Variational Emissions Inversion
The emissions inversions were performed using a variational algorithm, in which a cost function is minimized to find the optimal combination between a priori (background) and observational data. The cost function is minimized with respect to the increment $\Delta\textbf{x}$, which when added to the background state $\textbf{x}^{\rm{b}}$ yields the analysis $^{\rm{a}}=\textbf{x}^{\rm{b}}+\Delta\textbf{x}$ x. In our case, the cost function $J$ is given by
@@
where $\textbf{d}=\textbf{y}-H(\textbf{x}^{\mathrm{b}})$ is the difference between the observations y and the background in observation space. $H$ is the (nonlinear) observation operator that maps the model state into observation space, while H denotes its linearization. The first term in the right hand side of Equation (10) quantifies the deviation from the background state and is weighted by the inverse of the background error covariance B. Similarly, the second term in the right hand side of Eq. (10) quantifies the deviation from the observations and is weighted by the inverse pf the observation error covariance R.
@@
where $\textbf{B}_{cc}^{1/2}$ and $\textbf{B}_{ee}^{1/2}$ are the square roots of $\textbf{B}_{cc}$ and $\textbf{B}_{ee}$, respectively. The cost function can then be expressed as
@@
The emission inversion then consists of minimizing this cost function with respect to the control vector $\boldsymbol{\chi}$, followed by the transformation of the optimal control vector to the emissions (or concentration) increment given by Equation (13b).
@@
where $\{\delta\textbf{e}_{i}\}$ and $\{\delta\textbf{c}_{i}\}$ are emissions and concentration perturbation ensembles, respectively, $\mathrm{diag}(\cdot)$ is the diagonalization operator, and $\mathcal{L}^{1/2}$ is the square root of the localization matrix. Here, we have chosen the same localization matrix for both the emissions and concentrations, which was defined to be isotropic with a half-width at half-maximum of 120 km with the shape prescribed by the fifth-order function of Gaspari and Cohn (1999).
@@ Appendix B Components of the U-Net
The different components that comprise the U-Net used for EnsAI are described in this appendix. In this appendix, all references to graphical representations of the U-Net components refer to Fig. 3.
Residual blocks (He et al., 2016) (represented by the blue right arrows) are used throughout the network and consist of two sequential constitutional layers with a residual connection. Each constitutional layer is comprised by (in sequential order) a $3\times 3$ two-dimensional convolution, a group normalization (Wu and He, 2018), and a Sigmoid Linear Unit (SiLU) (Hendrycks and Gimpel, 2016) activation function.
@@ Appendix C Details of Correlation Length Decomposition
[…]
@@
[…]
@@
Appendix S2 Additional Plots for the Direct Comparison of Covariances
Figure S2: Statistics for the standard deviation of surface ammonia $\sigma_{c}$ taken over the combined evaluation time period of July 1-7 and August 1-7. The left panel shows the mean value of $\sigma_{c}^{\rm{GM}}$ over this time, while the middle and right panels show the $RMSE(\sigma_{c})$ values for the EnsAI ensemble and static covariance, respectively.
Figure S3: $RMSE$ values for the horizontal correlation length modes $L_{ec}[m]$ for $m\leq 2$ for the emissions/surface concentration cross-correlation over the combined time period of July 1-7 and August 1-7 for the EnsAI ensemble (left column) and the static covariance (right column). A narrow band of grid points around the lateral boundaries of the model grid is excluded as the full mode expansion of $L_{ec}$ cannot be computed at these points.

### v2

@@ 8.4 Comparison Results for Spatial Correlations
To conclude this section, we briefly examine the higher correlation length modes with $m>2$. Figure 15 shows the $ACC$ and $RMSE$ values of $L_{ec}[m]/L_{ec}[0]$ for $1\leq m\leq 11$ for June 2015. As seen in left panel of Fig. 15, the $ACC$ value for the EnsAI ensemble is above 0.85 for $m=1$ and decreases with $m$ to a value 0.56 by $m=10$. In contrast, the $ACC$ values for the static covariance show little change with wavenumber and is below 0.06 for all wavenumbers. From the right plot, we can see that the $RMSE$ values for both the EnsAI ensemble and static covariance decrease with wavenumber, as all three covariances models typically have less features on smaller scales (the decreasing values of $L_{ec}[m]$ with increasing $m$ that was shown in the example in Fig. 13 is a typical feature of the horizontal correlations). The $RMSE$ values for EnsAI are smaller than those for the static covariance for all wavenumbers, but the difference between the two decrease with increasing values of $m$. ACCACC and $RMSE$ values for the other months of the year look similar to those from June and can be found in Figures S9 and S10 of the Supplement.
Figure 15: ACCACC (left panel) and RMSERMSE (right panel) values for the horizontal correlation length modes for the emissions/surface concentration correlation $L_{ec}[m]$ as a function of wavenumber $m$ for EnsAI (orange) and the static covariance (green) for June 2015. The wavelength of the mode is displayed on the top axis.
@@ 9 Inversions
Sections 8.3 and 8.4 showed the results of direct comparisons between the covariances to determine how well EnsAI replicated the GEM-MACH ensemble perturbations. For some applications, such as model uncertainty quantification, the determination of the model’s error covariance is the end goal. However, for inversions/assimilation, the objective is to produce an analysis (as well as forecasts initialized by the analysis in the case of assimilation). In this section, we compare emissions inversion results made using the different model error covariances.
@@
As the inversions are only sensitive to the model error covariances at times and locations near the available observations, not all differences between the error covariances will show in the comparison between the increments and will be dependent on the particular observations used. As such, pairing the comparison of inversion results with the direct comparison between the covariances in the previous section provides a robust examination between the different error covariances.
The ammonia emissions increments produced using the inversion scheme described in Section 2 using the GEM-MACH ensemble, EnsAI ensemble, and static error covariances (denoted by $\Delta\textbf{e}^{\rm{GM}}$, $\Delta\textbf{e}^{\rm{EnsAI}}$, and $\Delta\textbf{e}^{\rm{static}}$, respectively) can be found in Figures S11, S12, and S13 of the Supplement. We highlight the differences between these increments in Figure 16, which shows the deviation between the GEM-MACH ensemble and static covariance increments $|\Delta\textbf{e}^{\rm{GM}}-\Delta\textbf{e}^{\rm{static}}|$ minus the deviation between the GEM-MACH ensemble and EnsAI ensemble increments $|\Delta\textbf{e}^{\rm{GM}}-\Delta\textbf{e}^{\rm{EnsAI}}|$. In this figure, locations where the EnsAI increment is closer to the GEM-MACH increment are shown in red and locations where the static covariance increment is closer to the GEM-MACH increment are shown in blue. As seen in Fig. 16, in all months more locations have the EnsAI increment closer to the GEM-MACH increment than the static covariance increment is (although the difference is not as large for January). For most months, there are many locations where the EnsAI increment closer to the GEM-MACH increment by between 25 and 50 $\rm{mg}\;\rm{s}^{-1}\;\rm{km}^{-2}$ (see Fig. 1 for comparison to the annual mean emissions). While some locations in the winter have the static covariance increment is closer to the GEM-MACH increment by between 10 and 20 $\rm{mg}\;\rm{s}^{-1}\;\rm{km}^{-2}$, most locations where the static covariance increment is closer to the GEM-MACH increment are closer by less than 10 $\rm{mg}\;\rm{s}^{-1}\;\rm{km}^{-2}$.
Figure 16: Differences between ammonia emissions increments $\Delta\textbf{e}$ from inversions using different background error covariances. Plots show the deviation between the GEM-MACH ensemble and static covariance increments $|\Delta\textbf{e}^{\rm{GM}}-\Delta\textbf{e}^{\rm{static}}|$ minus the deviation between the GEM-MACH ensemble and EnsAI ensemble increments $|\Delta\textbf{e}^{\rm{GM}}-\Delta\textbf{e}^{\rm{EnsAI}}|$.
Figure 17 shows the root mean square error of the ammonia emissions increment $RMSE(\Delta\textbf{e})$. Uncertainties were added to each curve using the analysis error covariance for the inversion, but are not easily visible in Fig. 17 as the uncertainty on the $RMSE(\Delta\textbf{e})$ values are small. The $RMSE(\Delta\textbf{e})$ for the increments produced using the static covariances are between 33% and 70% larger than the $RMSE$ values for EnsAI, with the exception of the January increment where $RMSE(\Delta\textbf{e})$ for the static covariances is only slightly higher than for EnsAI. Therefore, overall, the increments produced using the EnsAI ensemble are much closer to the increments produced by the static covariances (except for the January where the EnsAI increment is only slightly closer).
Figure 17: The root mean square error of the ammonia emissions increment $RMSE(\Delta\textbf{e})$ between EnsAI and GEM-MACH (orange curve) and the static covariance and GEM-MACH (green curve).
@@ 10 Conclusions
While application to emissions inversion systems is a substantial development by itself, the longer term goal is to expand EnsAI to more atmospheric chemical constituents to use in chemical data assimilation. Although EnsAI was able to reproduce most the time-dependent features of the ammonia ensemble with only surface fields as input, more vertical levels may be needed if applied to other species with longer atmospheric life times.
@@ Appendix A Ensemble-Variational Emissions Inversion
Emissions inversions were performed using a variational algorithm, in which a cost function is minimized to find the optimal combination between a priori (background) and observational information. The cost function is minimized with respect to the increment $\Delta\textbf{x}$, which when added to the background state $\textbf{x}^{\rm{b}}$ yields the analysis $\textbf{x}^{\rm{a}}=\textbf{x}^{\rm{b}}+\Delta\textbf{x}$. In our case, the cost function $J$ is given by
@@
where $\textbf{d}=\textbf{y}-H(\textbf{x}^{\mathrm{b}})$ is the difference between the observations y and the background in observation space. $H$ is the (nonlinear) observation operator that maps the model state into observation space, while H denotes its linearization. The first term in the right hand side of Equation (10) quantifies the deviation from the background state and is weighted by the inverse of the background error covariance B. Similarly, the second term in the right hand side of Eq. (10) quantifies the deviation from the observations and is weighted by the inverse of the observation error covariance R.
@@
where $\textbf{B}_{cc}^{1/2}$ and $\textbf{B}_{ee}^{1/2}$ are the square roots of $\textbf{B}_{cc}$ and $\textbf{B}_{ee}$, respectively. The cost function can then be expressed in terms of the control vector as
@@
The emissions inversion then consists of minimizing this cost function with respect to the control vector $\boldsymbol{\chi}$, followed by the transformation of the optimal control vector to the emissions (or concentration) increment given by Equation (13b).
@@
where $\{\delta\textbf{e}_{i}\}$ and $\{\delta\textbf{c}_{i}\}$ are the emissions and concentration perturbation ensembles, respectively, $\mathrm{diag}(\cdot)$ is the diagonalization operator, and $\mathcal{L}^{1/2}$ is the square root of the localization matrix. Here, we have chosen the same localization matrix for both the emissions and concentrations, which was defined to be isotropic with a half-width at half-maximum of 120 km with the shape prescribed by the fifth-order function of Gaspari and Cohn (1999).
@@ Appendix B Components of the U-Net
The different components that comprise the U-Net used for EnsAI are described in this appendix. All references to graphical representations of the U-Net components refer to Fig. 3.
Residual blocks (He et al., 2016) (represented by the blue right arrows) are used throughout the network and consist of two sequential convolutional layers with a residual connection. Each convolutional layer is comprised by (in sequential order) a $3\times 3$ two-dimensional convolution, a group normalization (Wu and He, 2018), and a Sigmoid Linear Unit (SiLU) (Hendrycks and Gimpel, 2016) activation function.
@@ Appendix C Details of Correlation Length Decomposition
$L[0]=\sqrt{\frac{2}{\pi}}\int r^{2}\varrho(r,\theta)drd\theta,$ (16a)
$\begin{split}L[1]&=\sqrt{\frac{2}{\pi}}\int(\cos(\phi)+i\sin(\phi))r^{2}\varrho(r,\theta)drd\theta\\ &=\sqrt{\frac{2}{\pi}}\int(x+iy)\varrho(x,y)dxdy,\end{split}$ (16b)
$\begin{split}L[2]&=\sqrt{\frac{2}{\pi}}\int(\cos(2\phi)+i\sin(2\phi))r^{2}\varrho(r,\theta)drd\theta\\ &=\sqrt{\frac{2}{\pi}}\int(\cos^{2}(\phi)-\sin^{2}(\phi)+i\sin(\phi)\cos(\phi))r^{2}\varrho(r,\theta)drd\theta\\ &=\sqrt{\frac{2}{\pi}}\int(x^{2}-y^{2}+ixy)\frac{\varrho(x,y)}{\sqrt{x^{2}+y^{2}}}dxdy,\end{split}$ (16c)
@@
$\begin{split}L[0]&=\frac{1}{\sqrt{2}\pi^{3/2}L_{G}^{2}}\int_{0}^{\infty}r^{2}\exp\left(-\frac{r^{2}}{2L_{G}^{2}}\right)dr\int_{0}^{2\pi}d\theta\\ &=\frac{\sqrt{2}}{\sqrt{\pi}L_{G}^{2}}\int_{0}^{\infty}r^{2}\exp\left(-\frac{r^{2}}{2L_{G}^{2}}\right)dr\\ &=L_{G},\end{split}$ (19)
@@
Figure S2: Time means of the standard deviation of surface ammonia $\sigma_{c}$ and the ammonia emissions/surface concentration isotropic correlation length $L_{ec}[0]$ taken over 2015.
Figure S3: $RMSE(\sigma_{c})$ values for the surface ammonia error standard deviation $\sigma_{c}$ for EnsAI (left column) and the static error covariances (right column) for December-January-February (DJF), March-April-May (MAM), June-July-August (JJA), and September-October-November (SON) of 2015.
Figure S4: ACCACC values for $L_{cc}[0]$, $L_{cc}[1]/L_{cc}[0]$, and $L_{cc}[2]/L_{cc}[0]$ over the model domain for each month in 2015.
Figure S5: Same as Figure S3 but for $RMSE(L_{ec}[0])$.
Figure S6: Same as Figure S3 but for $RMSE(L_{ec}[1]/L_{ec}[0])$.
Figure S7: Same as Figure S3 but for $RMSE(L_{ec}[2]/L_{ec}[0])$.
Figure S8: RMSERMSE values for $L_{cc}[0]$, $L_{cc}[1]/L_{cc}[q0]$, and $L_{cc}[2]/L_{cc}[0]$ over the model domain for each month in 2015.
Figure S9: ACCACC values for the horizontal correlation length modes for the emissions/surface concentration correlation $L_{ec}[m]$ as a function of wavenumber $m$ for EnsAI (orange) and the static covariance (green). The wavelength of the mode is displayed on the top axis.
Figure S10: Same as Figure S9 but for $RMSE$.
Figure S11: Ammonia emissions increments from inversions using the GEM-MACH ensemble to form the background error covariances.
Figure S12: Same as Figure S11 but using the EnsAI ensemble to form the background error covariances.
Figure S13: Same as Figure S11 but using the static covariances as the background error covariances.